I am having problems with Algebra/Trigonometry sticking in my brain. I did this problem 3x+4= -8
-4= 4
_____________________
x = -4
What I don't understand is what happens to the 3???
My next problem is 2(x+7)=5x+14 I don't know where to start??

Answer by Edwin McCravy(20086)

You skipped a couple of steps

We started with this

3x+4 = -8

Then we wanted to get the +4 off of the left side so we
add -4 to both sides like this

3x+4 = -8
  -4 = -4

Then we draw a line:

3x+4 = -8
  -4 = -4
----------

Then we bring down the 3x, we don't write anything under
the +4 and -4 because they cancel.  We bring down the
= sign and we combine -8 and -4 and get -12.

3x+4 = -8
  -4 = -4
----------
3x   = -12

You skipped that step.  Next we want to get rid of the 3.
That is a multiplied 3, so we will need to divide both
sides by 3, to do that we put a fraction bar under each side
and put a 3 as a denominator, like this

3x+4 = -8
  -4 = -4
----------
3x   = -12

3x     -12 
———— = ————— 
 3       3

You also skipped that step. Now the 3's cancel and just leave 
an x on the left, and we divide the -12 by 3 and get -4

3x+4 = -8
  -4 = -4
----------
3x   = -12

 3x     -12 
———— = ————— 
 3       3

   x = -4  

----------------------------------

We start with this:

2(x+7) = 5x+14

Use the distributive property to remove the parentheses on the left:

2(x+7) =  5x+14
 2x+14 =  5x+14

 Next we add -14 to both sides and draw a line:

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 --------------

Now we bring down the 2x, the +14 and -14 cancel so we don't
write anything there, then we bring down the = sign, then the
5x and we don't write anything at the end because that +14 and
-14 also cancel:

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 --------------
 2x    =  5x

Next we add -5x to both sides to get the 5x off the right side.

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 -------------
 2x    =  5x
-5x    = -5x

We draw a line 

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 -------------
 2x    =  5x
-5x    = -5x
--------------  

We combine the 2x and the -5x and get -3x, we bring
down the = sign, now ordinarily we wouldn't write anything under 
the right side since 5x and -5x cancel, but we MUST NEVER
LEAVE a side of the equation EMPTY!  So we must put a 0 there:

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 --------------
 2x    =  5x
-5x    = -5x
---------------
-3x    =  0

Next we want to get rid of the -3.
That is a multiplied -3, so we will need to divide both
sides by -3.  To do that we put a fraction bar under each side
and put a -3 as a denominator, like this

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 -------------
 2x    =  5x
-5x    = -5x
--------------
-3x    =  0

-3x      0 
———— = ————— 
-3      -3

Then the -3's cancel on the left. On the right
we are simply dividing 0 by -3 and that just
gives us 0.  So the whole thing worked out is:

2(x+7) =  5x+14
 2x+14 =  5x+14
   -14      -14
 -------------
 2x    =  5x
-5x    = -5x
--------------
-3x    =  0

-3x      0 
———— = ————— 
-3      -3

   x = 0  

Edwin